OpenStudy (anonymous):

what is {cscx/(1+cscx)}-{cscx/(1-cscx)}

OpenStudy (anonymous):

Are you trying to simplify it?

OpenStudy (anonymous):

yes haha forgot to put that

OpenStudy (anonymous):

I can give you a computer generated solution, which shows it being equal to sin(x)^2-1

Parth (parthkohli):

Oh, why do we need computers.

OpenStudy (anonymous):

that actually helps. cause now i can work backwards

Parth (parthkohli):

\[\dfrac{\csc(x)}{1 - \csc(x)} - \dfrac{\csc(x)}{1 + \csc(x)} = \dfrac{\csc(x) + \csc^2(x)}{1 + \csc^2 (x)} - \dfrac{\csc(x) - \csc^2(x)}{1 + \csc^2(x)}\]